can u please help me with this equation?. How many milliliters of a salt solution with a density of 1.8 g/mL are needed to provide 396 g of a salt solution?
Answer by stanbon(75887)

can u please help me with this equation?. How many milliliters of a salt solution with a density of 1.8 g/mL are needed to provide 396 g of a salt solution?
------------------
x/396g = 1mL/(1.8g)
x = 396/1.8 = 220 mL
---------------------------
Cheers,
Stan H.
===========================